OpenStudy (anonymous):

twice the number of reds exceeded three times The number of blues by 8. The ratio of the reds the the sum of the reds and blues was 5 to 7. How many were red and how many were blue?

OpenStudy (anonymous):

Im not sure if this is right: 2(r)+8=3(b)

OpenStudy (anonymous):

that should be fine

OpenStudy (anonymous):

also you need the equation R /R+B = (5/7)

OpenStudy (anonymous):

Take these two and solve

OpenStudy (anonymous):

how would you solve it

OpenStudy (anonymous):

ive tried

OpenStudy (anonymous):

7r=5r+5b, 2r=5b, so r=5b/2 and we already have 2r+8=3b, , put the value of r-5b/2 in second equation